OpenStudy (anonymous):

Using the properties of equality, find the value of x in the equation below. 5(3x + 6) = 8

OpenStudy (anonymous):

Equality is a very powerful tool. When you have an equation (an expression with equality), you can do anything to it as long as you do it to both sides.

OpenStudy (anonymous):

So you can move things around, like \[5(3x + 6) = 8\] to \[5(3x + 6) -8 = 0\] for example.

OpenStudy (anonymous):

In your case you want to find x, So what you want to do is to move things around x apart so that x ends up on one side and everything else if possible is on the other.

OpenStudy (anonymous):

answer it as a fraction with the problem i gave u

OpenStudy (anonymous):

A good first step would be to divide both sides by five, giving you \[\frac{ 5(3x + 6) }{ 5 } = \frac{ 8 }{ 5 }\]

OpenStudy (anonymous):

Or \[(3x + 6) = \frac{ 8 }{ 5 }\]

OpenStudy (anonymous):

And you can then get rid of the brackets on the left side. (brackets only mean something if they separate the inside from the outside, when there's nothing outside the brackets you can just ignore them.) giving you: \[3x + 6 = \frac{ 8 }{ 5 }\]

OpenStudy (anonymous):

Can you think of the next step?

OpenStudy (anonymous):

no i need the answer

OpenStudy (anonymous):

So you want to have only x on the left, and a number on the right. That's what solving for x means. So we moved the five, and now the expression looks more like what we want. But we have this plus sign, and a six, and that's very distracting. if we move the six to the right, we'll only have a 3x on the left, and that looks a lot better. Can you do that?

OpenStudy (anonymous):

no i want to not move them i want the asnwer as a fraction

OpenStudy (anonymous):

Well, how are you going to find the answer without moving the numbers?